29-14 only

29-14 only
media%2F522%2F522f4864-edea-4093-8a38-e0
Answer #1

(i) o,NNH,(aq)^ o,NNH-(aq)+ H(aq) (fast equilibrium) 2) o,NNH(aN,olg)+a(slow) (3) H (aq)+OH-(aq)-> H,ov) (fast)

For my convenience I denote O₂NNHʂ with N and O₂ NN H- with I (intermediate). Then r = k[I]. From the equilibrium one gets K= k1/k-1= [I][H+]/[N]. Solve for [I] and insert in to rate law: r= k2 k1 [N]/ ( k-1 [H+] ) and by comparison: k= k1 k2/k-1